ELECTRONIC SPECTROSCOPY OF JET-COOLED 2-METHOXYNAPHTHALENE

\underline{T. TROXLER}, Departement für Chemie und Biochemie, Universität Bern, Freiestrasse 3, CH-3012 Bern, Switzerland ^{a}.

The S_{1}-S_{0} electronic transition of jet-cooled 2-methoxynaphthalene (2MXN) has been investigated by use of laser-induced fluorescence spectroscopy. The fluorescence excitation spectrum revealed two electronic origin transitions, separated by 660 cm^{−1}. They are caused by the presence of a cis and a trans conformation of the methoxy group with respect to the naphthalene frame. Dispersed fluorescence spectra of 25 major transitions in the excitation spectrum of the cis conformer were recorded, employing a liquid nitrogen cooled charge-coupled device (CCD) detector. Atypical, highly structured emission indicates that extensive vibrational mixing in the S_{1} electronic state of cis-2MXN occurs, caused by large normal coordinate (Duschinsky) rotation. Experimentally determined frequencies are compared to the predicted normal mode values of an ab-initio calculation at the HF/6-31G^{∗∗} level. The spectrum, molecular structure, harmonic force field and vibrational coupling will be discussed.

aSupport by the Swiss National Science Foundation is gratefully acknowledged.